i need your answer about this..10+x=14
Answer by Edwin McCravy(8909)



need your answer about this..10+x=14

       10 + x = 14

The terms on the left are 10 and x. 

There is one term on the right, 14.

The idea is to get rid of all terms 

on the left that do not contain x.

So we get rid of the 10.  To get rid 

of the 10 we add its oppposite, -10, 

to both sides of the equation.  That 

is, we write -10 under the 10 and 

also under the 14,  then draw a line 

underneath like this:

       10 + x = 14
      -10      -10 
    ----------------

Then the 10 and the -10 cancel on the 

left and we write nothing underneath. 

We bring down the x and the equal sign. 

Then we combine the 14 and the -10 and 

get 4, so we end up with:

       10 + x = 14
      -10      -10 
    ----------------
            x = 4

So the solution is 4.

Edwin